# Large Language Model-Brained GUI Agents: A Survey

### v9

@@ 9.3 Evaluation Platforms
WebWalkerQA [368] Web 2025 Benchmarks the capacity of LLMs to handle deep, structured, and realistic web-based navigation and reasoning tasks. 680 high-quality QA pairs. Task Success Rate, Efficiency Score. Text Match, Action Match. https://github.com/Alibaba-NLP/WebWalker
@@ 9.4 Web Agent Benchmarks
Evaluating GUI agents in web environments necessitates benchmarks that capture the complexities and nuances of web-based tasks. Over the years, several benchmarks have been developed, each contributing unique perspectives and challenges to advance the field. We first provide an overview of these benchmarks in Tables XXIII, XXIV and XXV.
@@
Beyond the challenges of multimodality, understanding agents’ resilience to environmental distractions is crucial. EnvDistraction [362] introduces a benchmark that evaluates the faithfulness of multimodal GUI agents under non-malicious distractions, such as pop-ups and recommendations. The study demonstrates that even advanced agents are prone to such distractions, revealing vulnerabilities that necessitate robust multimodal perception for reliable automation.
@@
Addressing the automation of enterprise software tasks, WorkArena [360] offers a benchmark emphasizing tasks commonly performed within the ServiceNow platform. With 19,912 unique instances across 33 tasks, it highlights the significant performance gap between current state-of-the-art agents and human capabilities in enterprise UI automation, setting a trajectory for future innovation.
BrowserGym [367] builds ecosystem designed for web agent research. It unifies various benchmarks like MiniWoB(++) [140], WebArena [350], and WorkArena [360] under a single framework, addressing the issue of fragmentation in web agent evaluation. By leveraging standardized observation and action spaces, it enables consistent and reproducible experiments. BrowserGym’s extensible architecture make it a vital tool for developing and testing GUI-driven agents powered by LLMs, significantly accelerating innovation in web automation research.
@@
TABLE XXVI: Overview of mobile GUI agent benchmarks (Part I).
@@
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
AndroidEnv [240] Android 2021 Provides an open-source platform based on the Android ecosystem with over 100 tasks across approximately 30 apps, focusing on reinforcement learning for various Android interactions. 100+ tasks NA NA https://github.com/google-deepmind/android_env
@@
PIXELHELP [138] Android 2020 Includes a corpus of natural language instructions paired with UI actions across four task categories, aiding in grounding language to UI interactions. 187 multi-step instructions Step Success Rate Element Match, Action Match https://github.com/google-research/google-research/tree/master/seq2act
@@
Mobile-Env [370] Android 2024 Comprehensive toolkit for Android GUI benchmarks to enable controlled evaluations of real-world app interactions. 224 tasks Task Success Rate, Step Success Rate Text Match, Element Match, Image Match, State Information https://github.com/X-LANCE/Mobile-Env
@@
B-MOCA [371] Android 2024 Benchmarks mobile device control agents on realistic tasks, incorporating UI layout and language randomization to evaluate generalization capabilities. 131 tasks Task Success Rate Element Match, State Information https://b-moca.github.io/
@@
AndroidWorld [372] Android 2024 Offers a dynamic Android environment, allowing for diverse natural language instruction testing. 116 tasks Task Success Rate State Information https://github.com/google-research/android_world
@@
Mobile-Eval [152] Android 2024 Benchmark based on mainstream Android apps, and designed to test common mobile interactions. 30 instructions Task Success Rate, Step Success Rate, Efficiency Score Text Match, Path Length https://github.com/X-PLUG/MobileAgent
@@
DroidTask [150] Android 2024 Android Task Automation benchmark supports exploration and task recording in real apps with corresponding GUI action traces. 158 tasks Step Success Rate, Task Success Rate Element Match, Action Match https://github.com/MobileLLM/AutoDroid
@@
AITW [310] Android 2023 A large-scale dataset, which is partly inspired by PIXELHELP, covering diverse Android interactions. 715,142 episodes Task Success Rate, Step Success Rate Action Match https://github.com/google-research/google-research/tree/master/android_in_the_wild
@@
AndroidArena [373] Android 2024 Focuses on daily cross-app and constrained tasks within the Android ecosystem, providing single-app and multi-app interaction scenarios. 221 tasks Task Success Rate, Step Success Rate, Efficiency Score Action Match, Path Length https://github.com/AndroidArenaAgent/AndroidArena
TABLE XXVII: Overview of mobile GUI agent benchmarks (Part II).
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
ANDROIDLAB [317] Android 2024 Provides a structured evaluation framework with 138 tasks across nine apps, supporting both text-only and multimodal agent evaluations on Android. 138 tasks Task Success Rate, Step Success Rate, Efficiency Score Element Match, Image Match https://github.com/THUDM/Android-Lab
@@
GTArena [379] Mobile applications 2024 Introduces a Transition Tuple for GUI defects, enabling large-scale defect dataset creation and reproducible, end-to-end automated testing. 10,000+ GUI display and GUI interactions Task Success Rate, Step Success Rate. Text Match, Element Match, Action Match. https://github.com/ZJU-ACES-ISE/ChatUITest
@@
A3 [380] Mobile Android 2025 Introduces a novel business-level LLM-based evaluation process, significantly reducing human labor and coding expertise requirements. 201 tasks across 21 widely used apps. Task Success Rate. Element Match, Action Match. https://yuxiangchai.github.io/Android-Agent-Arena/
@@
LlamaTouch [377] Mobile Android 2024 Enables faithful and scalable evaluations for mobile UI task automation by matching task execution traces against annotated essential states 496 tasks covering 57 unique Android applications Task Success Rate, Step Success Rate, Efficiency Score Text Match, Action Match, State Information Match https://github.com/LlamaTouch/LlamaTouch
@@
MobileAgentBench [378] Mobile Android 2024 Provides a fully autonomous evaluation process on real Android devices and flexibility in judging success conditions across multiple paths to completion 100 tasks across 10 open-source Android applications Task Success Rate, Efficiency Score, Latency, Token Cost State Information (UI State Matching) https://mobileagentbench.github.io/
@@
Mobile-Bench [374] Android 2024 Supports both UI and API-based actions in multi-app scenarios, testing agents on single and multi-task structures with a checkpoint-based evaluation approach. 832 entries (200+ tasks) Task Success Rate, Step Success Rate, Efficiency Score Action Match, Path Length https://github.com/XiaoMi/MobileBench
@@
Mobile Safety Bench [375] Android 2024 Prioritizes safety evaluation in mobile control tasks, with distinct tasks focused on helpfulness, privacy, and legal compliance. 100 tasks Task Success Rate, Risk Mitigation Success Action Match with Safety Considered, Element Match, State Information https://mobilesafetybench.github.io/
@@
SPA-BENCH [376] Android 2024 Extensive evaluation framework supporting single-app and cross-app tasks in English and Chinese, providing a plug-and-play structure for diverse task scenarios. 340 tasks Task Success Rate, Step Success Rate, Efficiency Score Action Match, State Information, Time Spent, API Cost https://spa-bench.github.io
@@
SPHINX [381] Android 2025 Provides a fully automated benchmarking suite and introduces a multi-dimensional evaluation framework. 284 common tasks. Task Success Rate, Efficiency Score, Completion under Policy, Turn Success Rate. Text Match, Image Match, Element Match, Action Match. /
@@ 9.5 Mobile Agent Benchmarks
Evaluating GUI agents on mobile platforms presents unique challenges due to the diversity of interactions and the complexity of mobile applications. Several benchmarks have been developed to address these challenges, each contributing to the advancement of mobile agent evaluation. We first provide an analysis for these mobile benchmarks in Tables XXVI and XXVII.
@@
Building upon the need for systematic evaluation, ANDROIDLAB [317] establishes a comprehensive framework for Android-based autonomous agents. It introduces both an action space and operational modes that support consistent evaluations for text-only and multimodal models. By providing XML and SoM operation modes, ANDROIDLAB allows LLMs and LMMs to simulate real-world interactions in equivalent environments. The benchmark includes 138 tasks across nine apps, encompassing typical Android functionalities, and evaluates agents using metrics such as Success Rate and Reversed Redundancy.
To further challenge agents in handling both API and UI operations, Mobile-Bench [374] offers an innovative approach by combining these elements within a realistic Android environment. Its multi-app setup and three distinct task categories test agents’ capabilities in handling simple and complex mobile interactions, pushing beyond traditional single-app scenarios. The evaluation leverages CheckPoint metrics, assessing agents at each key action step, providing insights into planning and decision-making skills.
Emphasizing safety in mobile device control, MobileSafetyBench [375] provides a structured evaluation framework that prioritizes both helpfulness and safety. It rigorously tests agents across common mobile tasks within an Android emulator, focusing on layered risk assessment, including legal compliance and privacy. A distinctive feature is its indirect prompt injection test to probe agent robustness. The evaluation ensures agents are scored on practical success while managing risks, advancing research in LLM reliability and secure autonomous device control.
@@
Focusing on efficient and user-friendly evaluation, MobileAgentBench [378] presents a benchmark tailored for agents on Android devices. It offers a fully autonomous testing process, leveraging final UI state matching and real-time app event tracking. With 100 tasks across 10 open-source Android applications categorized by difficulty, it accommodates multiple paths to success, enhancing reliability and applicability. Comprehensive metrics, including task success rate, efficiency, latency, and token cost, provide insights into agent performance.
Complementing these efforts, LlamaTouch [377] introduces a benchmark and testbed for mobile UI task automation in real-world Android environments. Emphasizing essential state annotation, it enables precise evaluation of tasks regardless of execution path variability or dynamic UI elements. With 496 tasks spanning 57 unique applications, LlamaTouch demonstrates scalability and fidelity through advanced matching techniques, integrating pixel-level screenshots and textual screen hierarchies, reducing false negatives and supporting diverse task complexities.
Zhao et al., introduce GTArena [379], a formalized framework and benchmark designed to advance autonomous GUI testing agents. GTArena provides a standardized evaluation environment tailored for multimodal large language models. Central to its design is the novel Transition Tuple data structure, which systematically captures and analyzes GUI defects. The benchmark assesses three core tasks—test intention generation, task execution, and defect detection—using a diverse dataset comprising real-world, artificially injected, and synthetic defects, establishing GTArena as a pioneering benchmark for GUI testing agents.
@@
TABLE XXVIII: Overview of computer GUI agent benchmarks.
@@
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
OSWorld [357] Linux, Windows, macOS 2024 Scalable, real computer environment for multimodal agents, supporting task setup, execution-based evaluation, and interactive learning across Ubuntu, Windows, and macOS. 369 Ubuntu tasks, 43 Windows tasks Task Success Rate Execution-based State Information (such as internal file interpretation, permission management) https://os-world.github.io/
@@
Windows Agent Arena [382] Windows 2024 Adaptation of OSWorld focusing exclusively on the Windows OS with diverse multi-step tasks, enabling agents to use a wide range of applications and tools. 154 tasks Task Success Rate Same as OSWorld, scalable with cloud parallelization https://microsoft.github.io/WindowsAgentArena
@@
OmniACT [358] MacOS, Linux, Windows 2024 Assesses agents’ capability to generate executable programs for computer tasks across desktop and web applications in various OS environments, prioritizing multimodal challenges. 9,802 data points Task Success Rate, Step Success Rate Action Match https://huggingface.co/datasets/Writer/omniact
@@
VideoGUI [359] Windows 2024 Focuses on visual-centric tasks from instructional videos, emphasizing planning and action precision in applications like Adobe Photoshop and Premiere Pro. 178 tasks, 463 subtasks Task Success Rate State Information, Action Match https://showlab.github.io/videogui
@@
Spider2-V [383] Linux 2024 Benchmarks agents across data science and engineering workflows in authentic enterprise software environments, covering tasks from data ingestion to visualization. 494 tasks Task Success Rate Action Match, State Information https://spider2-v.github.io
@@
Act2Cap [292] Windows 2024 Emphasizes GUI action narration using cursor-based prompts in video format, covering a variety of GUI interactions like clicks, typing, and dragging. 4,189 samples Step Success Rate Element Match https://showlab.github.io/GUI-Narrator
@@
OFFICEBENCH [384] Linux 2024 Tests cross-application automation in office workflows with complex multi-step tasks across applications like Word and Excel, assessing operational integration in realistic scenarios. 300 tasks Task Success Rate Action match, Text Match, State Information https://github.com/zlwang-cs/OfficeBench
@@
AssistGUI [103] Windows Platform 2024 The first benchmark focused on task-oriented desktop GUI automation 100 tasks from 9 popular applications Task Success Rate, Efficiency Score Element Match, Action Match https://showlab.github.io/assistgui/
@@
WorldGUI [386] Windows platform 2025 First GUI benchmark designed to evaluate dynamic GUI interactions by incorporating various initial states. 315 total tasks from 10 Windows applications Task Success Rate, Efficiency Score Image Match, Element Match, Action Match /

### v11

@@ 9.3 Evaluation Platforms
WebWalkerQA [421] Web 2025 Yes Benchmarks the capacity of LLMs to handle deep, structured, and realistic web-based navigation and reasoning tasks. 680 high-quality QA pairs. Task Success Rate, Efficiency Score. Text Match, Action Match. https://github.com/Alibaba-NLP/WebWalker
WebGames [422] Web 2025 Yes A comprehensive benchmark designed to evaluate the capabilities of general-purpose web-browsing AI agents through 50+ interactive challenges. It uniquely provides a hermetic testing environment with verifiable ground-truth solutions. 50+ challenges Task Success Rate Action Match https://github.com/convergence-ai/webgames
TABLE XXXV: Overview of web GUI agent benchmarks (Part IV).
Benchmark Platform Year Live Highlight Data Size Metric Measurement Link
SafeArena [423] Web 2025 The first benchmark specifically designed to evaluate the deliberate misuse of web agents by testing their ability to complete both safe and harmful tasks. 500 tasks Task Success Rate, Completion under Policy, Risk Ratio Text Match, State Information https://safearena.github.io
WABER [424] Web 2025 Yes Introduces two new evaluation metrics—Efficiency and Reliability—that go beyond standard success rate measurements 655 tasks Task Success Rate, Efficiency Score Action Match, State Information https://github.com/SumanKNath/WABER
Online-Mind2Web [425] Web 2025 Yes A real-world online evaluation benchmark designed to reflect actual user interactions with live web interfaces 300 tasks from 136 websites Task Success Rate, Efficiency Score Image Match, Action Match, State Information, LLM-as-a-Judge Evaluation https://github.com/OSU-NLP-Group/Online-Mind2Web
AgentDAM [426] Web 2025 Yes The first benchmark to evaluate privacy leakage risks in multimodal, realistic web environments using agentic models 246 human-annotated test cases Task Success Rate, Risk Ratio Action Match, Text Match https://github.com/facebookresearch/ai-agent-privacy
AgentRewardBench [427] Web 2025 No The first benchmark to rigorously evaluate LLM-based judges against human expert annotations across multiple web agent tasks 1,302 trajectories, 351 tasks Task Success Rate, Completion under Policy Image Match, Element/State Match https://agent-reward-bench.github.io
RealWebAssist [428] Web 2025 No The first benchmark to evaluate long-horizon web assistance using real-world users’ sequential instructions expressed in natural and often ambiguous language 1,885 instructions Task Success Rate, Step Success Rate, Efficiency Score Action Match https://scai.cs.jhu.edu/projects/RealWebAssist/
REAL [429] Web 2025 Yes Fully deterministic, high-fidelity replicas of real-world websites (e.g., Airbnb, Amazon, Gmail), enabling safe, reproducible, and configurable testing for multi-turn GUI-based agents 112 tasks across 11 deterministic websites Task Success Rate Text Match, Action Match, State Information Match https://github.com/agi-inc/agisdk
BEARCUBS [430] Web 2025 Yes Emphasizes interaction with live web pages and includes multimodal tasks (e.g., video, audio, 3D) that cannot be solved by text-only methods, addressing limitations of prior benchmarks relying on static or simulated environments 111 questions Task Success Rate, Efficiency Score Text Match, Action Match https://bear-cubs.github.io
@@ 9.4 Web Agent Benchmarks
Evaluating GUI agents in web environments necessitates benchmarks that capture the complexities and nuances of web-based tasks. Over the years, several benchmarks have been developed, each contributing unique perspectives and challenges to advance the field. We first provide an overview of these benchmarks in Tables XXXII, XXXIII, XXXIV and XXXV.
@@
Beyond the challenges of multimodality, understanding agents’ resilience to environmental distractions is crucial. EnvDistraction [416] introduces a benchmark that evaluates the faithfulness of multimodal GUI agents under n-malicious distractions, such as pop-ups and recommendations. The study demonstrates that even advanced agents are prone to such distractions, revealing vulnerabilities that necessitate robust multimodal perception for reliable automation.
@@
Addressing the automation of enterprise software tasks, WorkArena [414] offers a benchmark emphasizing tasks commonly performed within the Service w platform. With 19,912 unique instances across 33 tasks, it highlights the significant performance gap between current state-of-the-art agents and human capabilities in enterprise UI automation, setting a trajectory for future in vation.
BrowserGym [420] builds ecosystem designed for web agent research. It unifies various benchmarks like MiniWoB(++) [145], WebArena [406], and WorkArena [414] under a single framework, addressing the issue of fragmentation in web agent evaluation. By leveraging standardized observation and action spaces, it enables consistent and reproducible experiments. BrowserGym’s extensible architecture make it a vital tool for developing and testing GUI-driven agents powered by LLMs, significantly accelerating in vation in web automation research.
@@
TABLE XXXVI: Overview of mobile GUI agent benchmarks (Part I).
@@
Benchmark Platform Year Live Highlight Data Size Metric Measurement Link
@@
AndroidEnv [263] Android 2021 Yes Provides an open-source platform based on the Android ecosystem with over 100 tasks across approximately 30 apps, focusing on reinforcement learning for various Android interactions. 100+ tasks NA NA https://github.com/google-deepmind/android_env
@@
PIXELHELP [143] Android 2020 No Includes a corpus of natural language instructions paired with UI actions across four task categories, aiding in grounding language to UI interactions. 187 multi-step instructions Step Success Rate Element Match, Action Match https://github.com/google-research/google-research/tree/master/seq2act
@@
Mobile-Env [431] Android 2024 Yes Comprehensive toolkit for Android GUI benchmarks to enable controlled evaluations of real-world app interactions. 224 tasks Task Success Rate, Step Success Rate Text Match, Element Match, Image Match, State Information https://github.com/X-LANCE/Mobile-Env
@@
B-MOCA [432] Android 2024 Yes Benchmarks mobile device control agents on realistic tasks, incorporating UI layout and language randomization to evaluate generalization capabilities. 131 tasks Task Success Rate Element Match, State Information https://b-moca.github.io/
@@
AndroidWorld [433] Android 2024 Yes Offers a dynamic Android environment, allowing for diverse natural language instruction testing. 116 tasks Task Success Rate State Information https://github.com/google-research/android_world
@@
Mobile-Eval [157] Android 2024 Yes Benchmark based on mainstream Android apps, and designed to test common mobile interactions. 30 instructions Task Success Rate, Step Success Rate, Efficiency Score Text Match, Path Length https://github.com/X-PLUG/MobileAgent
@@
DroidTask [155] Android 2024 Yes Android Task Automation benchmark supports exploration and task recording in real apps with corresponding GUI action traces. 158 tasks Step Success Rate, Task Success Rate Element Match, Action Match https://github.com/MobileLLM/AutoDroid
@@
AITW [352] Android 2023 No A large-scale dataset, which is partly inspired by PIXELHELP, covering diverse Android interactions. 715,142 episodes Task Success Rate, Step Success Rate Action Match https://github.com/google-research/google-research/tree/master/android_in_the_wild
@@
AndroidArena [434] Android 2024 Yes Focuses on daily cross-app and constrained tasks within the Android ecosystem, providing single-app and multi-app interaction scenarios. 221 tasks Task Success Rate, Step Success Rate, Efficiency Score Action Match, Path Length https://github.com/AndroidArenaAgent/AndroidArena
@@
ANDROIDLAB [357] Android 2024 Yes Provides a structured evaluation framework with 138 tasks across nine apps, supporting both text-only and multimodal agent evaluations on Android. 138 tasks Task Success Rate, Step Success Rate, Efficiency Score Element Match, Image Match https://github.com/THUDM/Android-Lab
@@
GTArena [440] Mobile applications 2024 No Introduces a Transition Tuple for GUI defects, enabling large-scale defect dataset creation and reproducible, end-to-end automated testing. 10,000+ GUI display and GUI interactions Task Success Rate, Step Success Rate Text Match, Element Match, Action Match https://github.com/ZJU-ACES-ISE/ChatUITest
TABLE XXXVII: Overview of mobile GUI agent benchmarks (Part II).
Benchmark Platform Year Live Highlight Data Size Metric Measurement Link
@@
A3 [441] Mobile Android 2025 Yes Introduces a novel business-level LLM-based evaluation process, significantly reducing human labor and coding expertise requirements. 201 tasks across 21 widely used apps. Task Success Rate. Element Match, Action Match. https://yuxiangchai.github.io/Android-Agent-Arena/
@@
LlamaTouch [438] Mobile Android 2024 Yes Enables faithful and scalable evaluations for mobile UI task automation by matching task execution traces against annotated essential states 496 tasks covering 57 unique Android applications Task Success Rate, Step Success Rate, Efficiency Score Text Match, Action Match, State Information Match https://github.com/LlamaTouch/LlamaTouch
@@
MobileAgentBench [439] Mobile Android 2024 Yes Provides a fully autonomous evaluation process on real Android devices and flexibility in judging success conditions across multiple paths to completion 100 tasks across 10 open-source Android applications Task Success Rate, Efficiency Score, Latency, Token Cost State Information (UI State Matching) https://mobileagentbench.github.io/
@@
Mobile-Bench [435] Android 2024 Yes Supports both UI and API-based actions in multi-app scenarios, testing agents on single and multi-task structures with a checkpoint-based evaluation approach. 832 entries (200+ tasks) Task Success Rate, Step Success Rate, Efficiency Score Action Match, Path Length https://github.com/XiaoMi/MobileBench
@@
Mobile Safety Bench [436] Android 2024 Yes Prioritizes safety evaluation in mobile control tasks, with distinct tasks focused on helpfulness, privacy, and legal compliance. 100 tasks Task Success Rate, Risk Mitigation Success Action Match with Safety Considered, Element Match, State Information https://mobilesafetybench.github.io/
@@
SPA-BENCH [437] Android 2024 Yes Extensive evaluation framework supporting single-app and cross-app tasks in English and Chinese, providing a plug-and-play structure for diverse task scenarios. 340 tasks Task Success Rate, Step Success Rate, Efficiency Score Action Match, State Information, Time Spent, API Cost https://spa-bench.github.io
@@
SPHINX [442] Android 2025 Yes Provides a fully automated benchmarking suite and introduces a multi-dimensional evaluation framework. 284 common tasks. Task Success Rate, Efficiency Score, Completion under Policy, Turn Success Rate. Text Match, Image Match, Element Match, Action Match. /
AEIA-MN [443] Mobile Android 2025 Yes Introduces the Active Environment Injection Attack (AEIA) framework that actively manipulates environmental elements (e.g., notifications) in mobile operating systems to mislead multimodal LLM-powered agents. 61 tasks (AndroidWorld) + 45 tasks (AppAgent) Task Success Rate, Risk Ratio, Efficiency Score Text Match, State Information, Action Match /
AutoEval [444] Mobile Android 2025 AutoEval Introduces a fully autonomous evaluation framework for mobile agents, eliminating the need for manual task reward signal definition and extensive evaluation code development. 93 tasks Task Success Rate Action Match, State Information /
LearnGUI [321] Mobile Android 2025 Yes The first benchmark to systematically study few-shot demonstration-based learning in mobile GUI agents, featuring both offline and online task environments Offline: 2,252 tasks with k-shot variants across 44 apps; Online: 101 interactive tasks across 20 apps Task Success Rate Action Match https://lgy0404.github.io/LearnAct
@@ 9.5 Mobile Agent Benchmarks
Evaluating GUI agents on mobile platforms presents unique challenges due to the diversity of interactions and the complexity of mobile applications. Several benchmarks have been developed to address these challenges, each contributing to the advancement of mobile agent evaluation. We first provide an analysis for these mobile benchmarks in Tables XXXVI and XXXVII.
@@
Building upon the need for systematic evaluation, ANDROIDLAB [357] establishes a comprehensive framework for Android-based auto mous agents. It introduces both an action space and operational modes that support consistent evaluations for text-only and multimodal models. By providing XML and SoM operation modes, ANDROIDLAB allows LLMs and LMMs to simulate real-world interactions in equivalent environments. The benchmark includes 138 tasks across nine apps, encompassing typical Android functionalities, and evaluates agents using metrics such as Success Rate and Reversed Redundancy.
To further challenge agents in handling both API and UI operations, Mobile-Bench [435] offers an in vative approach by combining these elements within a realistic Android environment. Its multi-app setup and three distinct task categories test agents’ capabilities in handling simple and complex mobile interactions, pushing beyond traditional single-app scenarios. The evaluation leverages CheckPoint metrics, assessing agents at each key action step, providing insights into planning and decision-making skills.
Emphasizing safety in mobile device control, MobileSafetyBench [436] provides a structured evaluation framework that prioritizes both helpfulness and safety. It rigorously tests agents across common mobile tasks within an Android emulator, focusing on layered risk assessment, including legal compliance and privacy. A distinctive feature is its indirect prompt injection test to probe agent robustness. The evaluation ensures agents are scored on practical success while managing risks, advancing research in LLM reliability and secure auto mous device control.
@@
Focusing on efficient and user-friendly evaluation, MobileAgentBench [439] presents a benchmark tailored for agents on Android devices. It offers a fully auto mous testing process, leveraging final UI state matching and real-time app event tracking. With 100 tasks across 10 open-source Android applications categorized by difficulty, it accommodates multiple paths to success, enhancing reliability and applicability. Comprehensive metrics, including task success rate, efficiency, latency, and token cost, provide insights into agent performance.
Complementing these efforts, LlamaTouch [438] introduces a benchmark and testbed for mobile UI task automation in real-world Android environments. Emphasizing essential state an tation, it enables precise evaluation of tasks regardless of execution path variability or dynamic UI elements. With 496 tasks spanning 57 unique applications, LlamaTouch demonstrates scalability and fidelity through advanced matching techniques, integrating pixel-level screenshots and textual screen hierarchies, reducing false negatives and supporting diverse task complexities.
Zhao et al., introduce GTArena [440], a formalized framework and benchmark designed to advance auto mous GUI testing agents. GTArena provides a standardized evaluation environment tailored for multimodal large language models. Central to its design is the vel Transition Tuple data structure, which systematically captures and analyzes GUI defects. The benchmark assesses three core tasks—test intention generation, task execution, and defect detection—using a diverse dataset comprising real-world, artificially injected, and synthetic defects, establishing GTArena as a pioneering benchmark for GUI testing agents.
@@
TABLE XXXVIII: Overview of computer GUI agent benchmarks.
@@
Benchmark Platform Year Live Highlight Data Size Metric Measurement Link
@@
OSWorld [413] Linux, Windows, macOS 2024 Yes Scalable, real computer environment for multimodal agents, supporting task setup, execution-based evaluation, and interactive learning across Ubuntu, Windows, and macOS. 369 Ubuntu tasks, 43 Windows tasks Task Success Rate Execution-based State Information (such as internal file interpretation, permission management) https://os-world.github.io/
@@
Windows Agent Arena [446] Windows 2024 Yes Adaptation of OSWorld focusing exclusively on the Windows OS with diverse multi-step tasks, enabling agents to use a wide range of applications and tools. 154 tasks Task Success Rate Same as OSWorld, scalable with cloud parallelization https://microsoft.github.io/WindowsAgentArena
@@
OmniACT [452] macOS, Linux, Windows 2024 No Assesses agents’ capability to generate executable programs for computer tasks across desktop and web applications in various OS environments, prioritizing multimodal challenges. 9,802 data points Task Success Rate, Step Success Rate Action Match https://huggingface.co/datasets/Writer/omniact
@@
VideoGUI [453] Windows 2024 No Focuses on visual-centric tasks from instructional videos, emphasizing planning and action precision in applications like Adobe Photoshop and Premiere Pro. 178 tasks, 463 subtasks Task Success Rate State Information, Action Match https://showlab.github.io/videogui
@@
Spider2-V [447] Linux 2024 Yes Benchmarks agents across data science and engineering workflows in authentic enterprise software environments, covering tasks from data ingestion to visualization. 494 tasks Task Success Rate Action Match, State Information https://spider2-v.github.io
@@
Act2Cap [323] Windows 2024 Yes Emphasizes GUI action narration using cursor-based prompts in video format, covering a variety of GUI interactions like clicks, typing, and dragging. 4,189 samples Step Success Rate Element Match https://showlab.github.io/GUI-Narrator
@@
OFFICEBENCH [448] Linux 2024 Yes Tests cross-application automation in office workflows with complex multi-step tasks across applications like Word and Excel, assessing operational integration in realistic scenarios. 300 tasks Task Success Rate Action Match, Text Match, State Information https://github.com/zlwang-cs/OfficeBench
@@
AssistGUI [108] Windows 2024 Yes The first benchmark focused on task-oriented desktop GUI automation 100 tasks from 9 popular applications Task Success Rate, Efficiency Score Element Match, Action Match https://showlab.github.io/assistgui/
@@
WorldGUI [449] Windows 2025 Yes First GUI benchmark designed to evaluate dynamic GUI interactions by incorporating various initial states. 315 total tasks from 10 Windows applications Task Success Rate, Efficiency Score Image Match, Element Match, Action Match /
UI-Vision [450] Desktop (Windows, Linux) 2025 No The first large-scale benchmark specifically designed for desktop GUI agents 8,227 query–label pairs in total Task Success Rate Action Match, Text Match https://uivision.github.io
Computer Agent Arena [451] Windows, Ubuntu, macOS 2025 Yes The first large-scale, open-ended evaluation platform for multimodal LLM-based agents in real desktop computing environments User-proposed tasks Task Success Rate Human evaluators https://arena.xlang.ai/